OpenStudy (anonymous):

Solve the system:9x + 2y = 2 21x + 6y = 4 Write your answers as an ordered pair (x,y)

OpenStudy (anonymous):

i got x=3 i need help finding y

OpenStudy (anonymous):

Plugin and solve

OpenStudy (anonymous):

i tried i got confused

OpenStudy (anonymous):

ill show you my work hold on one sec

OpenStudy (anonymous):

Do you understand algebra?

OpenStudy (anonymous):

yes i do to a point

OpenStudy (anonymous):

Just plugin your x value in the equation

OpenStudy (anonymous):

9(3)+2y=2 21(3)+6y=4 90+8y=6 subtract 90 8y= -84 divide 8 y=-10.5 and i dont think it is that

OpenStudy (anonymous):

Look up "pre-algebra" in google that might help you. You're not going to understand it without understanding the basics of algebra.

OpenStudy (anonymous):

(1/3,-1/2) To find x first,

OpenStudy (anonymous):

wait so that is the answer?

OpenStudy (anonymous):

9x+2y=2 21x+6y=4 I used the elimination, i multiply the first equation to -3 -3[9x+2y=2] -27x-6y=-6 Then, first equation plus second -27x-6y=-6 21x+6y=4 Then , -6x=-2 X= 1/3

OpenStudy (anonymous):

Yes x=1/3 and y= -1/2

OpenStudy (anonymous):

thank you!! :)